# Frustrated Quantum Magnetism on Complex Networks: What Sets the Total Spin

### v2

@@ Abstract
Consider equal antiferromagnetic Heisenberg interactions between qubits sitting at the nodes of a complex, nonbipartite network. We ask the question: How does the network topology determine the net magnetization of the ground state and to what extent is it tunable? By examining various network families with tunable properties, we demonstrate that (i) graph heterogeneity, i.e., spread in the number of neighbors, is essential for a nonzero total spin, and (ii) other than the average number of neighbors, the key structure governing the total spin is the presence of (disassortative) hubs, as opposed to the level of frustration. We also show how to construct simple networks where the magnetization can be tuned over its entire range across both abrupt and continuous transitions, which may be realizable on existing platforms. Our findings pose a number of fundamental questions and strongly motivate wider exploration of quantum many-body phenomena beyond regular lattices.
@@
Understanding how a network of pairwise interactions between dynamical units control their collective behavior is a unifying challenge for modern science. While condensed matter physics has traditionally focused on short-range interactions on a lattice, the multidisciplinary field of complex networks [1, 2, 3] has shown, over the past few decades, that the network topology itself can play a pivotal role [4]. In particular, structures not found in regular lattices, e.g., heterogeneity in the number of neighbors, the small-world effect [5], dissimilarity between adjacent nodes [6], and communities [7], can radically alter critical phenomena, disease spreading, synchronization, response to node failures, and other dynamical processes [4, 8, 9, 10, 11, 12].
@@
However, most of these findings are for classical systems. While there is a sizable literature on single-particle quantum dynamics [13, 14, 15, 16, 17] and quantum communication networks [16, 17], the exploration of quantum many-body networks is in its infancy [16]. A few isolated studies have found a strong impact of heterogeneity [18, 19] and an absence of small-world effect [20] in certain quantum phase transitions, faster information scrambling with sparse long-range bonds [21, 22], and an insensitivity to graph topology in the effect of local measurements [23]. However, there is no general understanding of how the network structure governs collective quantum phenomena, especially in frustrated systems. The question is not just of fundamental interest but also timely, as it would soon be possible to engineer arbitrary interaction graphs between qubits with long coherence times, most notably with trapped ions [24, 25, 26, 27] and superconducting circuits [28, 29, 30, 31, 32], among other platforms [33, 34, 35, 36, 37, 38].
@@
In the context of magnetism, networks can encode variable levels of geometric frustration, which is at the heart of exotic phases of matter such as spin liquids [39, 40]. Frustration arises whenever all bonds cannot be simultaneously satisfied, the simplest example being a triangle with antiferromagnetic bonds [41]. Studies of frustration in (undirected) complex networks have focused on social imbalance [42] and classical Ising spin-glass physics [8]. In contrast, we consider an antiferromagnetic Heisenberg model of qubits, $\hat{H}=\sum_{i,j}J_{i,j}\hat{\mathbf{S}}_{i}\cdot\hat{\mathbf{S}}_{j}$, where we usually take $J_{i,j}=J>0$ on all bonds to focus on the role of network topology as opposed to bond disorder. For a bipartite graph the ground state is obtained by oppositely aligning the two sublattices (of size $N_{A}$ and $N_{B}$), yielding a total spin $S_{\text{total}}=|N_{A}-N_{B}|/2$ [43]. On the other hand, for a nonbipartite lattice (e.g., Kagome) frustration can stabilize a quantum spin liquid composed of fluctuating singlets with $S_{\text{total}}=0$ but no conventional long-range order [44, 45, 46]. Here we explore what sets $S_{\text{total}}$ in a complex nonbipartite graph and to what extent it is tunable. Although far from characterizing the ground state, our findings reveal surprising ways in which network topology shapes collective behavior of such systems.
We study a variety of networks in which key structural properties such as heterogeneity, number of triangles, average degree (number of neighbors), degree-degree correlations, and average path length between two nodes can be varied independently. We present the $S_{\text{total}}$ distribution over thousands of graphs with $N=30$ qubits using matrix product states [47] in the ITensor library [48] with a singular-value cutoff of $10^{-14}$ and a maximum bond dimension of $5000$, for which almost all graphs converged to an integer total spin. The same statistical features are obtained from exact diagonalization for $N=20$. Our results show, in particular, that $S_{\text{total}}$ is not sensitive to the level of frustration. Instead, it can be tuned over almost its entire range by embedding “hubs” connected to many low-degree nodes in a wider network, which can be understood intuitively. These findings constrain the ground-state magnetic order and raise open questions on the wider role of frustration.
@@
Random graphs.—We start by examining (connected) random graphs of $N$ spins with $N_{e}$ bonds. Figure 1(a) shows the distribution of $S_{\text{total}}$ over an ensemble of such networks. We notice that $S_{\text{total}}$ is small compared to its upper bound, $S_{\text{max}}=N/2$, which one might expect considering we have antiferromagnetic coupling on all bonds. Second, as shown in Fig. 1(b), the average magnetization $\bar{S}_{\text{total}}$ falls monotonically with $N_{e}$ or, equivalently, the average degree $\bar{k}=\frac{2N_{e}}{N}$. This trend is consistent across all families of networks that we study, and it interpolates between known limiting cases: For $N_{e}^{\text{min}}=N-1$ one gets a random tree where a sublattice imbalance $|N_{A}-N_{B}|$ can give rise to a net magnetization, with $\bar{S}_{\text{total}}\approx 0.23\sqrt{N}$ (see Supplement [49]). Conversely, $N_{e}^{\text{max}}=N(N-1)/2$ gives an all-to-all graph, where $\hat{H}=\frac{1}{2}\hat{\mathbf{S}}_{\text{total}}^{2}$ up to a constant and the ground state has $S_{\text{total}}=0$.
@@
More importantly, Fig. 1(c) shows that $S_{\text{total}}$ is not sensitive to the frustration level of a graph but rather to its heterogeneity and assortativity [6]. We look at three measures of frustration [50]: (1) the frustration index $N_{f}$, which is the minimum number of bonds one needs to cut to make the graph bipartite [51, 50, 52], (2) the number of triangles, which is a more local measure, and (3) the average path length (number of hops) between two nodes, which controls how strongly a given spin can affect any other spin in the network, adding to the level of frustration. Surprisingly, we find that all three measures have weak or no correlation with $S_{\text{total}}$. Instead, $S_{\text{total}}$ is correlated with the amount of heterogeneity and assortativity: The former is quantified by $\Delta k$, the spread in the number of neighbors across the network [4], whereas the latter keeps track of degree-degree correlation, i.e., whether high-degree nodes connect to high-degree nodes $(A>0)$ or to low-degree nodes $(A<0)$ [6]. Figure 1(c) tells us that the magnetization is larger for more heterogeneous graphs where high-degree nodes link to low-degree nodes. Below we systematically vary these structural properties in turn to make sense of the correlations.
@@
Figure 1: (a) Bars show the distribution of total spin in the ground state of $\mathcal{N}=916$ random graphs with $N=30$ qubits and $N_{e}=45$ bonds with equal antiferromagnetic Heisenberg coupling. The dashed curve shows $S_{\text{total}}$ for Ising spins. (b) Ensemble average of $S_{\text{total}}$ vs average degree $\bar{k}=2N_{e}/N$ with $\mathcal{N}\approx 924$. (c) Correlation with various graph metrics. Points inside the gray band are statistically insignificant.
@@
The strongest correlation seen in Fig. 1(c) is with the total spin for (classical) Ising spins on the same network. The latter is given by an optimal bipartition (also known as the max-cut problem [51, 8]), where one divides the spins into two oppositely aligned groups $A$ and $B$ so as to maximize the number of $A$-$B$ bonds. As shown in Fig. 1, this bipolar state overestimates $S_{\text{total}}$ for the Heisenberg spins, especially at larger $\bar{k}$. Nonetheless, it is useful for interpreting some of the results.
@@
Figure 2: Sample graph (top), degree distribution (middle), and $S_{\text{total}}$ distribution (bottom) for random regular, random, and scale-free Barabási-Albert networks with $N=30$ and $\bar{k}\approx 4$. The power-law tail in (f) gives rise to hubs [circled in (c)] that want to polarize their neighborhoods, as shown in the inset.
@@
Importance of heterogeneity.—Suppressing heterogeneity results in the family of random regular graphs, where each spin has the same degree $k$. Figure 2(g) shows that for $k=3$ one almost always (more than $98\%$ of the cases) has $S_{\text{total}}=0$, and the same holds for $k>3$ (see Supplement [49]). This strong result shows that heterogeneity is essential for a nonzero total spin. Conversely, one can ramp up the heterogeneity by producing scale-free graphs [53] where the degree distribution $p_{k}$ follows a power law. A well known example is the Barabási-Albert model [54], for which $p_{k}\sim 1/k^{3}$. As shown in Fig. 2(i), these graphs exhibit much higher magnetizations. The enhancement can be understood as originating from “hubs,” i.e., nodes with a disproportionately high degree. Such hubs form a locally star structure, favoring a net spin imbalance [see Fig. 2(f) inset]. As we discuss below, this effect persists when hubs are embedded in a broader class of networks.
Figure 3: (a) Average total spin and (b) heterogeneity, given by the standard deviation $\Delta k$ of the degree distribution, for random graphs without short loops and $N=30$. (c) $\bar{S}_{\text{total}}$ for Barabási-Albert graphs with $N=30$, $\bar{k}=3.8$, and tunable number of triangles, measured by the mean clustering coefficient $C\in[0,1]$ [5]. Dashed curve shows $\bar{S}_{\text{total}}$ for Ising spins.
@@
Insensitivity to frustration.—An alternative argument for the unmagnetized nature of random regular graphs [Fig. 2(g)] could be that they have fewer short loops compared to random graphs [55, 56], so they appear locally treelike. Thus, $S_{\text{total}}$ is small due to lack of frustration, not because of homogeneity. To test this proposition, we directly vary the number of short loops in two different ways and see its effect on $S_{\text{total}}$: First, we uniformly sample from random graphs without loops of up to a given size $l$ [57]. We find that the total-spin distribution is almost unaffected by removing all triangles ($l=3$). As $l$ is increased further, $\bar{S}_{\text{total}}$ eventually drops to zero at some threshold $l=l_{c}$ [Fig. 3(a)]; however, $l_{c}$ depends strongly on the number of bonds $N_{e}$ and $\bar{S}_{\text{total}}$ does not appear to be sensitive as to whether $l$ is even or odd. Instead, it follows the variation in the degree distribution, which becomes progressively narrower with the removal of short loops [Fig. 3(b)]. Thus, we infer that $S_{\text{total}}$ is dictated by the heterogeneity and not by the frustration. Second, we follow Ref. [58] to augment the Barabási-Albert model with a triangle formation step, which allows one to tune the number of triangles without changing the degree distribution or the assortativity. As shown in Fig. 3(c), this yields a weak variation of $\bar{S}_{\text{total}}$ which is reproduced by the optimal bipartition. We also find that $\bar{S}_{\text{total}}$ does not vary appreciably with shorter path lengths in small-world graphs (see Supplement [49]).
@@
While this indifference to frustration may be counterintuitive, note the same holds on lattices, where both non-frustrated (e.g., square) and frustrated (e.g., Kagome) cases have $S_{\text{total}}=0$, albeit with very different magnetic order. Likewise, we expect frustration to be important in setting the order on general networks.
Figure 4: (a) Most disassortative and (b) most assortative networks with the same degree sequence. Node colors show a bipolar ground state of Ising spins. (b) Variation of $\bar{S}_{\text{total}}$ as the assortativity [6] is tuned for random graphs with $N=30$ and $\bar{k}=4$. Dashed curve shows the variation for Ising spins.
@@
Importance of disassortativity.—Not all heterogeneous networks have large magnetization. Another key metric is the assortativity coefficient $A\in[-1,1]$ [6]. Perfect assortativity ($A=1$) requires all nodes to have the same degree (for which $S_{\text{total}}\simeq 0$), whereas perfect disassortativity ($A=-1$) occurs in a star graph and, more generally, in complete bipartite graphs with sublattice imbalance [59] (where $S_{\text{total}}$ can be large). Crucially, one can tune $A$ without altering the degrees via rewiring pairs of bonds [60, 61] or by solving a linear optimization problem [59]. Thus, Figs. 4(a) and 4(b) show two graphs with very different values of $A$ but the same degrees (hence, the same heterogeneity). The first graph is the most disassortative and has $S_{\text{total}}=7$, whereas the latter is the most assortative and has $S_{\text{total}}=0$. The physical origin of this large disparity becomes clear if one examines the structure of these networks [60]: Strongly disassortative graphs have a section where few high-degree nodes (hubs) connect to many low-degree nodes in an approximately bipartite structure [Fig. 4(a)], which gives rise to a large spin imbalance. By contrast, in strongly assortative graphs high-degree nodes form a tightly connected group and low-degree nodes form a treelike structure [Fig. 4(b)], neither of which favors a large magnetization. Note that the large $S_{\text{total}}$ in the former case requires both the presence of hubs and disassortativity.
@@
Figure 4(c) shows that $\bar{S}_{\text{total}}$ falls sharply as we tune $A$ from its minimum to maximum value for random graphs. The same behavior is observed for Ising spins and for scale-free graphs (see Supplement [49]).
@@
Figure 5: (a) Graph with embedded hubs and (b) graph with no hubs generated by the growth model described in text. Node colors show the ground state for Ising spins and bond colors show ground-state correlations for our Heisenberg model. (c)–(f) Variation of the total-spin distribution with the structure parameter $p$ for $N=30$ and $\bar{k}=3.8$. Dashed curves are for Ising spins. (i) Corresponding variation of heterogeneity $\Delta k$, frustration index $N_{f}$, and assortativity $A$. (j) Average total spin as a function of $p$ and $\bar{k}=2m-m(m+1)/N$ for $N=30$.
@@
Tunable spin distribution.—Fortunately, we can use a simple model to grow networks with tunable heterogeneity and assortativity. It has a parameter $m\in\mathbb{Z}$ that sets the average degree and another parameter $p\in[0,1]$ that controls the structure. The protocol starts with a small clique of $m$ nodes, and in each step links a new node $i$ to $m$ existing nodes as follows: (1) Select an existing node $j$ at random, (2) with probability $p$ link $(i,j)$, (3) else link $i$ to a neighbor of $j$ with preferential attachment, i.e., with probability proportional to the neighbor’s degree. This is a variant of the “copy models” for the Web [62, 63, 64].
@@
As shown in Fig. 5(a), for $p=0$ we get strongly disassortative graphs where many low-degree “outer” nodes connect to a few “central” hubs. As $p$ is increased to $1$, the graphs become more homogeneous [Fig. 5(b)]. Thus, both the heterogeneity and the disassortativity fall monotonically with $p$, while the frustration level varies weakly, as in Fig. 5(i). Accordingly, we find that the total-spin distribution is peaked at a macroscopic value of $O(N)$ for $p=0$ and moves toward zero as $p$ goes to $1$ [Figs. 5(c)-5(h)]. In addition, $\bar{S}_{\text{total}}$ decreases with the average degree $\bar{k}\approx 2m$ [Fig. 5(j)] as in random graphs.
@@
In Fig. 5(c) the distribution is broad, covering almost the full range of $S_{\text{total}}$. Here, the maximum total spin occurs when all of the outer nodes link to the initial clique, producing a star configuration with $m$ giant central hubs [as in Fig. 6(a) without the outer ring]. For $N\gg m$ it is then energetically favorable to align the hub spins opposite to the outer spins, which gives $S_{\text{total}}=N/2-m$. However, during the growth stage a new node may connect to an outer node instead, which can then accumulate more bonds to become a mini hub, at the expense the initial clique which now has fewer bonds. In other words, the giant hubs can break into several smaller hubs, which then align opposite to the outer nodes [see Fig. 5(a)], reducing $S_{\text{total}}$. Note that the bonds between the hubs add frustration but do not change the total spin.

### v3

@@ Abstract
Consider equal antiferromagnetic Heisenberg interactions between qubits forming a complex, nonbipartite network. We ask the question: How does the network topology determine the net magnetization of the ground state and to what extent is it tunable? By examining over $75000$ networks of different families with tunable structural properties, we demonstrate that (i) heterogeneity in the number of neighbors is essential for a nonzero total spin, and (ii) apart from the number of neighbors, the key determinant is the presence of (disassortative) hubs, as opposed to the frustration level. In fact, one can vary the magnetization throughout its range by embedding such hubs. We also discuss simple, exactly solvable networks where such tunability leads to both abrupt and continuous transitions, with quantum effects giving rise to a diverging susceptibility. Our findings can be realized on emerging platforms and pose a number of fundamental questions, strongly motivating wider exploration of quantum many-body phenomena on complex networks.
@@
Understanding how a network of interactions between dynamical units control their collective behavior is a unifying challenge for modern science. While, traditionally, condensed matter physics has focused on short-range interactions on a lattice, the rapidly growing field of complex networks [1, 2, 3] has shown that the network topology itself can play a pivotal role [4]. In particular, structures absent in regular lattices, e.g., heterogeneity in the number of neighbors, the small-world effect [5], dissimilarity between adjacent nodes (called “dis-assortativity” [6]), and communities [7], can radically alter critical phenomena, disease spreading, synchronization, response to node failures, and other dynamical processes [4, 8, 9, 10, 11, 12, 13].
@@
However, most of these findings are for classical systems. While there is a sizable literature on single-particle quantum dynamics [14, 15, 16, 17, 18] and quantum communication networks [17, 18], the exploration of quantum many-body networks is in its infancy [18, 19, 20, 21, 22, 23, 24, 25, 26]. In particular, there is no general understanding of how network structure governs collective properties, especially of a frustrated quantum system. This question is timely, as it would soon be possible to design essentially arbitrary interaction graphs between qubits with long coherence times, most notably with trapped ions [27, 28, 29, 30] and superconducting circuits [31, 32, 33, 34, 35], among other setups [36, 37, 38, 39, 40, 41].
@@
In the context of magnetism, networks can encode variable levels of geometric frustration, which is at the heart of exotic phases of matter such as spin liquids [42, 43]. Frustration arises whenever all bonds cannot be simultaneously satisfied, the simplest example being a triangle with antiferromagnetic bonds [44]. Studies of frustration in network science have focused on social imbalance [45] and Ising spin-glass physics [8]. In contrast, here we consider an antiferromagnetic Heisenberg model of qubits, $\hat{H}=\sum_{\langle i,j\rangle}J_{i,j}\hat{\mathbf{S}}_{i}\cdot\hat{\mathbf{S}}_{j}$, where we generally take $J_{i,j}=1$ to focus on the role of network topology as opposed to bond disorder. For a bipartite graph, the two sublattices (of size $N_{A}$ and $N_{B}$) align oppositely, yielding a total spin $S_{\text{total}}=|N_{A}-N_{B}|/2$ [46]. On the other hand, for a nonbipartite lattice (e.g., Kagome) frustration can stabilize a quantum spin liquid composed of fluctuating singlets with $S_{\text{total}}=0$ but no conventional long-range order [47, 48, 49]. Here we explore what sets $S_{\text{total}}$ in a complex nonbipartite graph and to what extent it is tunable. Although far from fully characterizing the ground state, our findings reveal surprising ways in which network topology shapes collective behavior of such systems.
We examine more than $70$ network families of $30$ qubits with adjustable structural properties, each comprising an ensemble of $1000$ graphs, using state-of-the-art exact diagonalization based on the HPhi library [50, 51]. We find, surprisingly, that $S_{\rm total}$ is not sensitive to the level of frustration; however, it is fully tunable by embedding “hubs” connected to many low-degree nodes, allowing one to control spin states by varying the graph topology. We also discuss exactly solvable cases where quantum effects enhance the sensitivity to network parameters, producing a divergent susceptibility. Our results show that the role of heterogeneity in promoting spin alignment [4] extends to quantum systems, while uncovering the central importance of assortative correlations. They also constrain the magnetic ordering on complex networks and raise fundamental questions on the wider role of frustration.
@@
Random graphs—We start by examining (connected) Erdős–Rényi graphs of $N$ sites and $N_{e}$ bonds. Figure 1(a) shows the distribution of such an ensemble. We notice that $S_{\text{total}}$ is small compared to its upper bound, $N/2$, which one might expect as all bonds have antiferromagnetic coupling. Second, as shown in Fig. 1(b), the average magnetization $\bar{S}_{\text{total}}$ falls with $N_{e}$ or, equivalently, the average degree $\bar{k}=\frac{2N_{e}}{N}$. This trend is consistent across all network families and interpolates between known limiting cases: For $N_{e}^{\text{min}}=N-1$ one gets a random tree where a sublattice imbalance gives rise to a net magnetization with $\bar{S}_{\text{total}}\approx 0.23\sqrt{N}$; see Supplemental Material (SM) [52]. On the other hand, $N_{e}^{\text{max}}=N(N-1)/2$ gives an all-to-all graph, where $\hat{H}=\frac{1}{2}\hat{\mathbf{S}}_{\text{total}}^{2}$ up to a constant and the ground state has $S_{\text{total}}=0$.
@@
More importantly, Fig. 1(c) shows that $S_{\text{total}}$ is not sensitive to the frustration level of a graph but rather to its heterogeneity and assortativity [6]. We look at three measures of frustration [53]: (1) the frustration index, which is the minimum number of bonds one needs to cut to make the graph bipartite [54, 53, 55], (2) the number of triangles, which is a more local measure, and (3) the average path length (number of hops) between two nodes, which controls how strongly a given spin can affect any other spin in the network, adding to the level of frustration. Surprisingly, we find that all three measures have weak or no correlation with $S_{\text{total}}$. Instead, $S_{\text{total}}$ is correlated with the amount of heterogeneity and assortativity: The former is quantified by $\Delta k$, the spread in the number of neighbors [4], whereas the latter keeps track of degree-degree correlation, i.e., whether high-degree nodes connect to high-degree nodes $(A>0)$ or to low-degree nodes $(A<0)$ [6]. Figure 1(c) tells us that $S_{\text{total}}$ is larger for more heterogeneous graphs where high-degree nodes link to low-degree nodes. Below we vary these structural properties in turn to make sense of the correlations.
@@
Figure 1: (a) Distribution of total spin in the ground state of random graphs with $N=30$ qubits and $N_{e}=45$ bonds with equal antiferromagnetic Heisenberg coupling. Dashed curve shows the average magnetization, $S^{z}_{\text{total}}$, for Ising spins (max-cut bipartitions). (b) Ensemble average of $S_{\text{total}}$ vs average degree $\bar{k}=2N_{e}/N$. (c) Correlation with various graph metrics. Points inside the gray band are statistically insignificant.
@@
We also find a strong correlation ($\sim 0.6$) with the magnetization, $S^{z}_{\rm total}$, of Ising spins on the same graph. The latter is given by an optimal bipartition (known as the max-cut problem [8]), where one divides the spins into oppositely aligned groups $A$ and $B$ so as to maximize the number of AA-$B$ bonds. As shown in Fig. 1, such a bipolar state overestimates $S_{\text{total}}$ for the Heisenberg spins, especially at larger $\bar{k}$. Nonetheless, the major variations with structural properties are reflected in this purely topological metric. Note the solution to the max-cut problem is generally not unique, in which case we take the average $S^{z}_{\rm total}$ over the degenerate manifold.
@@
Figure 2: Sample graph (top), degree distribution (middle), and $S_{\text{total}}$ distribution (bottom) for random regular, random, and scale-free (Barabási-Albert) graphs with $N=30$ and $\bar{k}\approx 4$. The power-law tail in (f) gives rise to hubs [circled in (c)] that want to polarize their neighborhoods (inset). Dashed curves in (g-i) show the distribution of average $S^{z}_{\rm total}$ for Ising spins.
Importance of heterogeneity—Suppressing heterogeneity results in the family of random regular graphs, where each spin has the same degree $k$. Figure 2(g) shows that for $k=4$ one almost always has $S_{\text{total}}=0$, and the same holds for $k\neq 4$ (see SM [52]). This strong result shows that heterogeneity is essential for a nonzero total spin.
@@
Conversely, one can ramp up the heterogeneity by generating scale-free graphs [56] where the degree distribution $p_{k}$ follows a power law. A well known example is the Barabási-Albert model [57] for which $p_{k}\sim 1/k^{3}$. As shown in Fig. 2(i), these graphs exhibit much higher magnetizations. The enhancement can be understood as originating from “hubs,” i.e., nodes with a disproportionately high degree. Such hubs form a locally star structure, favoring a net spin imbalance [see Fig. 2(f) inset]. As we discuss below, this effect persists even when hubs are embedded in a broader class of networks.
Note that all three families in Fig. 2 have similar frustration indices ($\sim 10.1$, $10.5$, and $10.6$, respectively).
@@
Insensitivity to frustration—An alternative argument for the unmagnetized nature of random regular graphs [Fig. 2(g)] could be that they have fewer short loops compared to random graphs [58, 59], so they appear locally treelike. Thus, $S_{\text{total}}$ is small due to lack of local frustration, not homogeneity. To test this proposition, we vary the number of short loops in two different ways: First, we uniformly sample random graphs with a minimum loop size (girth) $l$ [60]. We find that the total-spin distribution is almost unaffected by removing all triangles. As $l$ is increased, $\bar{S}_{\text{total}}$ eventually drops to zero close to the maximum girth $l_{m}$ [Fig. 3(a)]; however, $l_{m}$ depends strongly on the number of bonds $N_{e}$ [61], and $\bar{S}_{\text{total}}$ does not appear to be sensitive to whether $l$ is even or odd. Instead, it follows the variation in the degree distribution, which becomes progressively narrower with the removal of short loops [Fig. 3(b)]. Thus, we infer that $S_{\text{total}}$ is dictated by the heterogeneity and not by the frustration. Second, we follow Ref. [62] to augment the Barabási-Albert model with a triangle formation step, which allows one to tune the number of triangles without affecting the degree distribution or assortativity. As shown in Fig. 3(c), this yields a weak variation of $\bar{S}_{\text{total}}$. We also find that $\bar{S}_{\text{total}}$ does not vary significantly with shorter path lengths in small-world networks (see End Matter).
Figure 3: (a) Average total spin and (b) heterogeneity, given by the standard deviation $\Delta k$ of the degree distribution, for random graphs without short loops and $N=30$. (c) $\bar{S}_{\text{total}}$ for Barabási-Albert graphs with $N=30$, $\bar{k}=3.8$, and tunable number of triangles, measured by the mean clustering coefficient $C\in[0,1]$ [5]. Dashed curve shows $\bar{S}^{z}_{\text{total}}$ for Ising spins.
@@
While this indifference to frustration may be counterintuitive, note the same holds on regular lattices, where both non-frustrated (square) and frustrated (Kagome) cases have $S_{\text{total}}=0$, albeit with distinct spin ordering. However, its extension to general networks is indeed surprising and demands a deeper explanation.
@@
Importance of disassortativity—Not all heterogeneous networks have large magnetization. Another key metric is the assortativity coefficient $A\in[-1,1]$ [6]. Perfect assortativity requires all nodes to have the same degree (for which $S_{\text{total}}\simeq 0$), whereas perfect disassortativity occurs in complete bipartite graphs with sublattice imbalance [63] (where $S_{\text{total}}$ can be large). Crucially, one can tune $A$ without altering the degrees via rewiring pairs of bonds [63, 64, 65]. Thus, Figs. 4(a) and 4(b) show two graphs with very different values of $A$ but the same degrees (hence, the same heterogeneity). The first graph is the most disassortative and has $S_{\text{total}}=7$, whereas the latter is the most assortative and has $S_{\text{total}}=0$. The physical origin of this large disparity becomes clear if one examines the structure of these networks [64]: Strongly disassortative graphs have a section where few high-degree nodes (hubs) connect to many low-degree nodes in an approximately bipartite structure [Fig. 4(a)], which gives rise to a large spin imbalance. By contrast, in strongly assortative graphs high-degree nodes form a tightly connected group and low-degree nodes form a treelike structure [Fig. 4(b)], neither of which favors a large magnetization. Note that the large $S_{\text{total}}$ in the former case requires both the presence of hubs and disassortativity.
@@
Figure 4(c) shows that $\bar{S}_{\text{total}}$ falls sharply as we tune $A$ from its minimum to maximum value for random graphs. The same is observed for scale-free graphs (End Matter).
Figure 4: (a) Most disassortative and (b) most assortative networks with the same set of degrees. Node colors show a bipolar ground state of Ising spins. (b) Variation of $\bar{S}_{\text{total}}$ as the assortativity [6] is tuned for random graphs with $N=30$ and $\bar{k}=4$. Dashed curve shows $\bar{S}^{z}_{\rm total}$ for Ising spins.
@@
Figure 5: (a) Graph with embedded hubs and (b) graph with no hub from the “copy model” described in text for $m=2$. Node colors show a ground state for Ising spins and bond colors show ground-state correlations for our Heisenberg model. (c)–(e) Variation of total-spin distribution with the structure parameter $p$ for $N=30$ and $m=2$. Red dots show the average excitation gap for a given $S_{\rm total}$. (f) Variation of heterogeneity $\Delta k$, frustration index $N_{f}$, and assortativity AA. (g) $\bar{S}_{\rm total}$ as a function of $p$ and $\bar{k}=2m-m(m+1)/N$ for $N=30$.
@@
Tunable spin distribution—We propose a simple model to grow networks with tunable heterogeneity and assortativity. It has a parameter $m\in\mathbb{Z}$ that sets the average degree and another parameter $p\in[0,1]$ that controls the structure. The protocol starts with a small clique of $m$ nodes, and in each step links a new node $i$ to $m$ existing nodes as follows: (1) Select a node $j$ at random, (2) with probability $p$ link $(i,j)$, (3) else link $i$ to a neighbor of $j$ with preferential attachment, i.e., with probability proportional to the neighbor’s degree. This is a variant of the “copy models” for the Web [66, 67, 68].
@@
As shown in Fig. 5(a), for $p=0$ we get strongly disassortative graphs where many low-degree “outer” nodes connect to a few “central” hubs. As $p$ is increased to $1$, the graphs become more homogeneous [Fig. 5(b)]. Thus, both heterogeneity and disassortativity fall off with $p$, while the frustration index is almost invariant [Fig. 5(f)]. Accordingly, we find the total-spin distribution is peaked at a macroscopic value of $O(N)$ for $p=0$ and moves toward zero as $p$ goes to $1$ [Figs. 5(c)-5(e)]. In fact, finite-size scaling for $m=2$ indicates that this variation with $p$ spans the full range of $\bar{S}_{\rm total}$ for $N\to\infty$ (see End Matter). Such large, tunable magnetization in the presence of frustration is not found on regular lattices.
@@
In Fig. 5(c) the maximum total spin occurs when all of the outer nodes link to the initial clique, producing $m$ giant hubs that align opposite to the outer spins, yielding $S_{\text{total}}=N/2-m$. However, if a new node connects to an outer node, the latter can then accumulate more bonds to become a hub at the expense the initial clique, reducing $S_{\text{total}}$ [Fig. 5(a)]. Note that the bonds between the hubs add frustration without changing the total spin.
In addition, we find that the excitation gap rises with $S_{\rm total}$, especially for $p=0$ [Figs. 5(c)-5(e)], showing that the hubs also lend more energetic stability. This may be anticipated as the gap for a ring is $O(1/N)$ [69], whereas that for a star graph is $O(1)$. In most cases, we find the gap is magnetic, involving changes in $S_{\rm total}$ by $\pm 1$.
Deterministic tuning of $S_{\text{total}}$ in a frustrated network—Motivated by these results, we explore the phase diagram of a simple, hubbed network shaped like a wheel. It consists of $N_{c}$ central spins with all-to-all coupling $J_{c}$, $N_{b}$ outer spins forming a ring with coupling $J_{b}$, and all “spokes” connecting them with coupling $J$. Figure 6(a) shows a sketch for $N_{c}=2$. One recovers distinct topologies as one of these couplings is reduced toward zero: For $J_{b}=0$ and $J_{c}\sim J$, it is a $p=0$ graph [Fig. 5(a)] with $S_{\rm total}=(N_{b}-N_{c})/2\gg 1$. For $J_{c}=0$, it is equivalent to a Heisenberg ring in a (quantized) magnetic field, which should undergo a transition from a saturated ferromagnet to a Luttinger liquid with decreasing total spin as $J_{b}/J$ is increased [70]. For $J=0$, the ring is disconnected from the center, yielding $S_{\rm total}=0$ (if $N_{b}$, $N_{c}$ are even). Thus, by varying the couplings one should be able to stabilize any total spin between 0 and $(N_{b}-N_{c})/2$.